mikewwe13:

The total amount of money in a savings account after t years is given by the function A=1000(1.023)^t . How could this function be rewritten to identify the monthly interest rate? What is the approximate monthly interest rate? Drag and drop the choices into the boxes to correctly complete the table. If a value does not match, do not drag it to the table. Function Monthly interest rate A = 1000(1 + 0.023)^12t A = 1000(1.023^12)^t/12 A = 1000(1.023^t/12)^12t 0.23% 0.19% 0.31%

sillybilly123:

Multiply out your options: \(A = 1000(1 + 0.023)^{12t} = 1000(1 + 0.023)^{12t} \) !! \(A = 1000(1.023^{12})^{t/12} = 1000(1 + 0.023)^{t}\) ??? \(A = 1000(1.023^{t/12})^{12t}= 1000(1.023)^{t^2}\) :/ They're quite meaningless. To decompound \(2.3 \%\) into a monthly rate \(i_m\), you can go down Voc's road by saying that \(i_m = \dfrac{2.3}{12} = 0.1916 \%\). Or you can see that with monthly compounding \((1 + i_m)^{12} = 1.023 \implies i _ m = 1.023^{1/12} - 1 \approx 0.1897 \%\) So you get an answer but what are you asking?!?! questioncove.com/study#/updates/59f8ac9f7fd0d0a32be04c00

mikewwe13:

i mean like this

mikewwe13:

The total amount of money in a savings account after t years is given by the function A=1000(1.023)^t . How could this function be rewritten to identify the monthly interest rate? What is the approximate monthly interest rate? Drag and drop the choices into the boxes to correctly complete the table. If a value does not match, do not drag it to the table. Function Monthly interest rate A = 1000(1 + 0.023)^12t 0.23% 0.19% A = 1000(1.023^12)^t/12 0.31% A = 1000(1.023^t/12)^12t

mikewwe13:

like that

mikewwe13:

place to where it's right

sillybilly123:

none of them match up IMHO